Solve the following quadratic equation by factoring:
4x^2 - 9x + 2 = 0

Answer by edjones(8007)

4x^2 - 9x + 2 = 0
4*2=8. Two factors of 8 whose sum is -9 are -8 and -1
4x^2-8x-x+2=0
4x(x-2)-(x-2)=0
(4x-1)(x-2)=0
4x=1 x=1/4, x=2
.
Ed